OpenStudy (anonymous):

can someone teach me implicit and explicit differentiation like I'm 5?

OpenStudy (across):

What is it that you don't understand?

OpenStudy (anonymous):

Core concepts. My book doesn't explain it very clearly. For my first problem I'm supposed to find the derivative by implicit differentiation.

OpenStudy (anonymous):

http://puu.sh/5cBFL.png

OpenStudy (across):

Well, you're given\[x^2y+xy^2=-2.\]Assuming that \(y\) is a function of \(x\), we differentiate implicitly as follows:\[2xy+x^2y'+y^2+2xyy'=0.\]Now it's all algebra; solve for \(y'\):\[y'=-\frac{y^2+2xy}{x^2+2xy}.\]Note that this is a differential equation.

OpenStudy (phi):

Implicit differentiation of the first term \[ \frac{d}{dx}(x^2y) \] we use the product rule d(uv)= u dv + v du \[ \frac{d}{dx}(x^2y)= x^2 \frac{d}{dx}y + y \frac{d}{dx} x^2 \\ x^2 \frac{dy}{dx} + y \ 2x \frac{d}{dx}x\\ x^2 \frac{dy}{dx} + 2xy \]

zepdrix (zepdrix):

Sam, do you understand why the \(\Large \dfrac{dy}{dx}\)'s are showing up in the problem?

OpenStudy (anonymous):

previously I've written differentiation as a prime, like Y' or f'(x), so not really zepdrix

zepdrix (zepdrix):

Primes are fine :) Just so long as you know where they're coming from. I'm gonna ramble on for a minute here lol. You don't have to read it if you don't want XD I just wanna try to simplify things a tad if I can. Take for example:\[\Large y=x^2\]You're probably used to simply writing the derivative operator with the prime on the left side of the equation. But what's actually happening is, we're taking the derivative of each side.\[\Large (y)'\quad=\quad (x^2)'\qquad\to\qquad y'\quad=2x\] The tricky part with implicit differentiation is remembering that stupid chain rule!! :P In the above example we could have applied the chain rule to the x^2, giving us:\[\Large y\quad=\quad x^2\qquad\to\qquad y'\quad=\quad 2x(x)'\qquad\to\qquad y'\quad=\quad 2x(1)\]But as you can see, the x' is insignificant. The derivative of x with respect to x is just 1.

zepdrix (zepdrix):

So if we take an expression like:\[\Large \left(xy\right)'\]Applying the product rule (similar to what phi showed you),\[\Large =\quad (x)'y+x(y)'\]But as we mentioned before, x' is simply 1,\[\Large =\quad (1)y+xy'\]

zepdrix (zepdrix):

With implicit differentiation, when you take the derivative, you'll get a bunch of y' showing up in different places. We need to group them up together so we can solve for y'.

zepdrix (zepdrix):

These guys already gave some nice examples :O You can let me know if you need to see another though.

OpenStudy (anonymous):

Phi can you elaborate those steps??

OpenStudy (phi):

which steps ?

OpenStudy (phi):

How about \[ \frac{d}{dx} x \] ? you know that the answer is 1. one way to write this is \[ \frac{d}{dx} x= \frac{dx}{dx} = 1 \] if instead we did a constant, we get \[ \frac{d}{dx} 2 = \frac{d\ 2}{dx} = 0 \] if that makes any sense. if y is a variable and depends on x we would expect that as x changes, y changes, and therefore the ration \[ \frac{dy}{dx} \] exists and means something

OpenStudy (phi):

*ratio

OpenStudy (anonymous):

I understand how to find basic derivatives, I'm just confused in the dy/dx

OpenStudy (phi):

what is \[ \frac{d}{dx} y^2 \] we use the chain rule d u^n = n u^(n-1) du \[ \frac{d}{dx} y^2 = 2 y \ \frac{d}{dx} y= 2y \frac{dy}{dx} = 2 y \ y'\]

OpenStudy (phi):

notice the difference between d/dx y^2 and d/dx x^2 \[ \frac{d}{dx} x^2 = 2x \frac{d}{dx} x = 2x \frac{dx}{dx}= 2x\]

OpenStudy (phi):

can you do \[ \frac{d}{dx} xy \]?

OpenStudy (anonymous):

I see the difference but I don't know how to apply it

OpenStudy (phi):

can you try d/dx (x y) ?